OpenStudy (anonymous):

please help me, I really can't understand this question. f(x)\[\int\limits_{x}^{0}\sqrt{dt/t^{3}-t ^{2}} dt \] what is f'(x)

OpenStudy (anonymous):

f(x)\[\int\limits_{0}^{x}\sqrt{dt/t^{3}-t ^{2}} dt \] what is f'(x)

OpenStudy (eyust707):

I believe you can solve this using the first fundamental thrm of calculus.

OpenStudy (anonymous):

I'm not sure how to apply the theorem with a variable as a limit

OpenStudy (anonymous):

what do I do with the dt in the numerator?

OpenStudy (anonymous):

What is that!!!?

OpenStudy (anonymous):

it is a typo is my guess

OpenStudy (anonymous):

the dt? definitely not a typo.

OpenStudy (anonymous):

figured it out, it wasn't a typo you're just supposed to replace the dt with 1. thanks, all.

OpenStudy (anonymous):

\[f(x)=\int_0^x\sqrt{\frac{dt}{t^3-t^2}}dt\] is definitely a typo